# When Should Selfish Miners Double-Spend?

### v3

@@ 3.5 Definitions and Notations
(Model parameters) The fraction of the total hashrate the attacker pool possesses is denoted as $\bm{\alpha}$. The remaining miners are considered to be honest and possess $\bm{\beta}=1-\alpha$ fraction of the hashrate. When a public fork exists, $\bm{\gamma}$ fraction of the honest miners prefer the adversarial branch.
Definition 3
($\bm{k}$-Confirmation rule) A block is said to be confirmed according to the $k$-confirmation rule if it is part of the longest public chain and $k$-deep, i.e., there are at least k1k-1 public blocks mined on top of it in the longest chain.
@@
    $\bm{L}$-stubborn mining strategy refers to the mining strategy for which the procedure is explicitly explained in Section 3.2.
@@
    Honest mining strategy refers to the mining strategy when the pool strictly follows the honest protocol and publishes a block as soon as it is mined. The strategy can be explicitly obtained by plugging $\bm{L=1}$ into the procedure explained in Section 3.2.
@@
    Selfish mining strategy refers to the mining strategy proposed by Eyal and Sirer [4]. The strategy can be explicitly obtained by plugging $\bm{L=2}$ into the procedure explained in Section 3.2. We also give an explicit summary of the attack in Appendix A as well as how $L$-stubborn mining generalizes the attack.
@@
    $\bm{S}$-stealth mining strategy refers to the mining strategy for which the procedure is explicitly explained in Section 3.4.
@@
(Revenue ratio) Revenue ratio of the pool refers to the fraction of the blocks created by the pool in the longest chain in the long run. In other words,
$\bm{\rho=\frac{\#blocks~created~by~pool}{\#total~blocks~in~longest~chain}}$ (7)
@@
    $\bm{\rho_{L}}$ is the revenue ratio of $L$-stubborn mining strategy, for which an explicit equation is provided in Theorem 1, Eq. (18). Here,
@@
      $\bm{\rho_{1}}$ is the revenue ratio when the pool employs the honest mining strategy, where
@@
      $\bm{\rho_{2}}$ is the revenue ratio of selfish mining strategy, where
@@
      $\bm{\rho_{\infty}}$ is the revenue ratio of equal fork stubborn strategy, where
      $\rho_{\infty}$ $=\frac{\alpha}{\beta}-\frac{(1-2\alpha)(1-\gamma)}{\beta\gamma}(1-\beta C((1-\gamma)\alpha\beta)).$ (10)
@@
    $\bm{\sigma_{S}}$ is the revenue ratio of $S$-stealth mining strategy, for which an explicit equation is provided in Theorem 5, Eq. (31).
@@
In the long run, for every coin issued by the protocol, the pool gets $\rho$ fraction of it. Note that, for a given $(\alpha,\gamma)$ pair, if a specific mining strategy by the attacker pool results in a revenue ratio higher than the revenue ratio of honest mining strategy, i.e., $\rho>\alpha$, the system is not incentive compatible for the $(\alpha,\gamma)$ pair.
@@
We note that the following are defined in Section 3.2: $L$, AA, $H$, AA-chain, $H$-chain, offset chain, offset event, matching, override, adopt actions. As it is not possible to define them without completely going through the algorithm of $L$-stubborn mining, we refer the reader to the first paragraph of Section 3.2 where they are defined.
@@ 4 Analysis of $L$-Stubborn Mining
We start by analysis of the revenue ratio of the attacker pool when it employs the $L$-stubborn mining strategy. Instead of building a complex Markov chain and going through the tedious analysis of steady-state distribution as it is done in [4], we can analyze the distinct possible paths an attack cycle can take. An attack cycle starts when an offset chain is defined, and ends when a new offset chain is defined. If during an attack cycle, the length of the $A$-chain reaches $L$, i.e., $A=L$, then we call the attack successful. On the other hand, if the length of the $H$-chain surpasses the length of the $A$-chain before $A$ reaches $L$, we call the attack cycle unsuccessful.
In Figure 1, we display the sample paths of $4$-stubborn mining, i.e., $L=4$. The $y$-axis and $x$-axis represent $A$ and $H$, i.e., the length of $A$-chain and $H$-chain, respectively. An attack starts from point $(0,0)$ and arrivals of adversarial (honest resp.) blocks are represented with light-green (orange resp.) arrows. A red arrow shows that the arrival of honest block has ended the cycle. A blue (red resp.) cross represents the endpoint of a successful (unsuccessful resp.) attack cycle. A dark-green arrow represents the arrival of an adversarial block that guarantees a successful attack cycle. The dark-green crosses represent the first time cycles are guaranteed to be successful and not ended yet. The cycles that pass through the dark-green crosses eventually end with an honest arrival in blue cross points (x,x1)(x,x-1) where $x>4$, i.e., when the difference of chain lengths reduces to $1$.
Note that there is a third dimension of Figure 1, which tracks the number of adversarial blocks in the prefix of $H$-chain since every arrival of honest block has two choices of chains to extend. Further, the dynamics of this dimension is Markovian since each honest arrival either increases the previous value to H1H-1 (with probability $\gamma$) or stays the same (with probability $1-\gamma$) and does not depend on anything else. This Markovian structure implies that we can separate the third dimension and treat it independently which simplifies our analysis, hence, we do not plot it.
Figure 1: Sample paths of $4$-stubborn mining.
@@
Proof:  The end of the unsuccessful attack cycle implies that the last block that arrives has to be honest and the adversary was never behind up until that the last block. As a result, the arrival sequence of honest and adversarial blocks creates a Dyck string with length $n<L$. There are $C_{n}$ number of such strings and together with the last honest block each has probability $\alpha^{n}\beta^{n+1}$. Hence, an attack cycle ends unsuccessfully with the honest block at height $n+1$ with probability
@@
Proof:  First note that, in the long run, the longest chain is essentially the offset chain that grows when an adopt or override happens. As a result, the revenue ratio of the attacker is the fraction of the blocks created by the attacker in the offset chain. As the offset event is a stopping time, we can simply focus on a single attack cycle and calculate the expected number of blocks created by the adversary that become part of the offset chain. Dividing this number to the expected growth of the offset chain in an attack cycle gives us the revenue ratio. In Lemma 1 and Lemma 2, we provide the probability for each possible outcome of an attack cycle. If an attack cycle is successful, we have an $A$-chain at length $L$ and the maximum height of the honest block is $H=m<L$ at some time within the attack cycle. Clearly, if Lm>1L-m>1, the adversary will continue the cycle until the difference goes down to $1$. Thus, given the attack is successful, using Wald’s identity provided in Section 2.3, the expected number of blocks that become part of the offset chain in this case is $L+\frac{(L-m-1)\alpha}{1-2\alpha}$ and all these blocks are adversarial. This explains $L_{s}$ in (18). On the other hand, if the attack cycle is unsuccessful, $i$ adversarial blocks still make it into the offset chain w.p. $P_{u}(n,i)$ and the total number of blocks that make it into the offset chain is $n+1$. As a result, these two cases are considered in the numerator and denominator of (18) as $L_{u,a}$ and $L_{u}$, respectively.  $\blacksquare$
@@ 4.1 Optimizing the Revenue Ratio
As we mentioned earlier, our $L$-stubborn mining strategy is a generalization and extension of the existing mining strategies of honest mining $(L=1)$, selfish mining [4] $(L=2)$, equal fork stubborn mining [6] $(L=\infty)$. As a result, for a given $(\alpha,\gamma)$ pair, a natural question is, what integer value of $L$ maximizes the revenue ratio $\rho_{L}$. To find the optimal $L$, one could evaluate $\rho_{L}$ in (18) and find the $L$ that provides the highest revenue ratio. As the calculations involve binomial coefficients and (18) involves fractions, it is not straightforward to find $L^{*}$. Instead, here, we provide a much simpler method to obtain the optimal $L$. To do so, we resort to the linearization method of the MDP problem [5] which is presented in Appendix A, and use it in a narrow-sense. Note that, every time a block is mined, the adversary faces a decision as to whether it should publish $A$-chain and end the cycle, or continue. Clearly, if AH>1A-H>1, the decision should be to continue the attack cycle as the adversary can gain additional advantage at no-cost in these situations. On the other hand, if $A=H+1$, the risk is imminent as the adversary loses the advantage of overriding the $H$-chain if the next arrival is honest. As a result, we can narrow down the decision problem to a specific case where $A=H+1$ and analyze the situation therein via the linearization method.
@@
Next, using the definitions above and intermediary results in Appendix B, we provide the following quasiconcavity result related to the $L$-stubborn mining.
@@
Theorem 2 essentially says that $\rho_{L}$ is quasiconcave for integer $L$. We relegate the proof of Theorem 2 to the Appendix B as it requires some intermediary definitions and results. These results also shed light on a simple algorithm that can be used to find $L^{*}$.
Next, we provide Algorithm 1 that finds the optimal $L^{*}$ of $L$-stubborn mining. For that, let $\rho_{1}$, $\rho_{2}$ and $\rho_{\infty}$ be defined as in Definition 5, which are the revenue ratios of honest mining, selfish mining [4], and equal-fork stubborn mining [6, 37], respectively. Note that each of these values can be obtained by (18), where for $\rho_{\infty}$, $L_{s}$ can be set to zero as the $H$-chain will catch the $A$-chain eventually, i.e., no attack ends successfully, but the revenue comes from matching. In Appendix B, we use the intermediary results to show that Algorithm 1 outputs $L^{*}$, which essentially solves the decision problem mentioned earlier.
@@ 4.2 $(k+1)$-Stubborn Mining and Double-Spending
Note that, even when $L=k+1$, a successful attack cycle does not immediately imply a successful double-spending. It only implies that the $A$-chain will make it into the offset chain eventually. Similarly, an unsuccessful attack cycle does not immediately imply that $tx$ will be confirmed. Next, we provide the probability of each event regarding $tx$ and $tx^{\prime}$ for $(k+1)$-stubborn mining with $k$-confirmation rule.
@@
Proof:  For Double-Spending to happen, one possibility is, we will observe $A>k$ and $H\leq k$ at some point during the cycle. The first time this happens during the cycle is the first time $A=k+1$ and $H=m$ which happens w.p. $P_{s}(k+1,m)$. Until this point, adversary matches every one of the $m$ honest blocks. From this point onwards, adversary does not need to match the honest blocks as the strategy will eventually publish the longer $A$-chain. Hence, $tx$ stays in the $H$-chain if none of the $m$ matching adversarial blocks make it into the prefix of the the honest block at height $H$, which happens with probability $(1-\gamma)^{m-\mathbbm{1}_{m=k}}$. Note that, if $H=k$, then it does not matter whether prefix chain is after the $k$th block or not, hence $\mathbbm{1}_{m=k}$ at the exponent. Another possibility for Double-Spending is, if $H$-chain catches $A$-chain with both having length $k$ and the next block is honest which has probability $P_{u}(k)$. In this situation, if $H$-chain switches its prefix on the last honest block from $tx$ to $tx^{\prime}$, which happens w.p. $(1-\gamma)^{k-1}\gamma$, then $tx$ gets discarded despite being confirmed earlier. For Service to happen, we will observe $L\geq H=A+1=m+1$, which ends the cycle with $H$-chain making it into the offset chain which happens w.p. $P_{u}(m)$. Further, $tx$ has to be in the prefix of the honest block at height $m+1$, which explains the term $(1-\gamma)^{m}$.  $\blacksquare$
Here, we remind the reader that $\rho_{k+1}$ is the revenue ratio of the pool when it employs the $(k+1)$-stubborn mining strategy and $\bar{L}=\sup\{L:\rho_{L}\geq\alpha\}$, i.e., the largest value of $L$ such that the revenue ratio of $L$-stubborn mining is higher than the honest mining strategy. Hence, due to the quasiconcavity proven in Theorem 2, under any $k$-confirmation rule with $k<\bar{L}$ is compromised by $(k+1)$-stubborn mining strategy since the adversary gains more revenue ratio by employing this strategy compared to the revenue ratio of the honest mining strategy, i.e., $\rho_{k+1}\geq\alpha$, and at the same time it can double-spend at no-cost. In other words, for a given $\alpha$ and $\gamma$, $k$-confirmation rule should be at least $k\geq\bar{L}$ to make sure that double-spending attack has loss in revenue ratio. Note that, if $\rho_{\infty}\geq\alpha$, no $k$-confirmation is safe as double-spending has zero-cost. From this perspective, the safety of a blockchain system fails well before adversarial proportion reaches $\alpha$. We investigate this phenomenon further in the numerical results section by providing $L^{*}$ and $\bar{L}$ for each $\alpha$ and $\gamma$.
@@ 4.3 Combined Revenue Ratio
Proof:  Note that the first term in (23) provides the sum of successful cycles with double-spending events. Each term in that sum corresponds to the first time $A=k+1$ and $H=m$, and none of the matched adversarial blocks make it into the prefix of the honest block as explained in the proof of Theorem 3. Given Wald’s identity and the fact that a cycle will continue until the difference goes down to $1$, all but last $k$ adversarial blocks will result in double-spending events. Thus, total extra block rewards adversary gets in a successful attack cycle in addition to the $(k+1)_{s}$ is provided as in (27), where the extra term is $\mathcal{R}$ times the expected number of honest blocks that are confirmed and later replaced in an attack cycle. An unsuccessful cycle can only result in double-spending if the event corresponding to the second term in (23) happens, where only one honest block is confirmed and replaced later, which explains the extra term in (28).  $\blacksquare$
@@
Next, we provide another approach that incorporates the profit coming from double-spending when considering the $(k+1)$-stubborn mining attack provided here. As mentioned in [3], the value of service provided might be below the fee paid. In this case, we assume that each successful cycle results in only single double-spending event, i.e., only the first block in the $H$-chain is attacked which contains $tx$. We further assume that the value of service provided by the merchant equals $V$ and adversary pays $F$ fee for the service with $F\geq V$, where both values are in block reward units. In this case, on average, the $(k+1)$-stubborn mining attack is more profitable than honest mining if

### v4

@@ 3.4 Definitions and Notations
(Model parameters) The fraction of the total hashrate the adversary possesses is denoted as $\bm{\alpha<0.5}$. The remaining miners are considered to be honest and possess $\bm{\beta=1-\alpha}$ fraction of the hashrate. When a public fork exists, $\bm{\gamma}$ fraction of the honest miners prefer the adversarial branch.
@@
    $\bm{L}$-stubborn mining strategy refers to the mining strategy for which the procedure is explicitly explained in Section 3.1 and formally defined in Appendix B.3.
@@
    Honest mining strategy refers to the mining strategy when the adversary strictly follows the honest protocol and releases a block as soon as it is mined. The strategy can be explicitly obtained by plugging $\bm{L=1}$ into the procedure explained in Section 3.1.
@@
    Selfish mining strategy refers to the mining strategy proposed by Eyal and Sirer [4]. The strategy can be explicitly obtained by plugging $\bm{L=2}$ into the procedure explained in Section 3.1. We also give an explicit summary of the attack in Appendix B as well as how $L$-stubborn mining generalizes the attack.
@@
    $\bm{S}$-stealth mining strategy refers to the mining strategy for which the procedure is explicitly explained in Section 3.3 and formally defined in Appendix B.3.
@@
(Revenue ratio) Revenue ratio of the adversary refers to the fraction of the blocks created by the adversary in the longest chain in the long run. In other words, let $N_{A}^{t}$ and $N_{H}^{t}$ denote the the number of adversarial and honest blocks in the longest chain at time $t$ respectively. Then,
$\rho=\lim_{t\to\infty}\frac{N_{A}^{t}}{N_{A}^{t}+N_{H}^{t}}.$ (1)
@@
    $\rho_{L}$ is the revenue ratio under the $L$-stubborn mining strategy, for which an explicit equation is provided in Theorem 1, equation (12). Here:
@@
      $\rho_{1}$ is the revenue ratio when the adversary employs the honest mining strategy, where
@@
      $\rho_{2}$ is the revenue ratio of selfish mining strategy, where
@@
      $\rho_{\infty}$ is the revenue ratio of equal fork stubborn strategy, where
      $\rho_{\infty}$ $=\frac{\alpha}{\beta}-\frac{(1-2\alpha)(1-\gamma)}{\beta\gamma}(1-\beta C((1-\gamma)\alpha\beta))$ (4)
      and $C(x)$ is given in (40).
@@
    $\sigma_{S}$ is the revenue ratio of $S$-stealth mining strategy, for which an explicit equation is provided in Theorem 5, equation (26).
@@
In the long run, for every coin issued by the protocol, the adversary gets $\rho$ fraction of it. Note that, for a given $(\alpha,\gamma)$ pair, if a specific mining strategy by the adversary results in a revenue ratio higher than the revenue ratio of honest mining strategy, i.e., $\rho>\alpha$, the system is not incentive compatible for the $(\alpha,\gamma)$ pair.
@@
We note that the following are defined in Section 3.1: $L$, AA, $H$, AA-chain, $H$-chain, offset chain, offset event, matching, override, adopt actions. As it is not possible to define them without completely going through the algorithm of $L$-stubborn mining, we refer the reader to Section 3.1 where they are defined.
@@ 4 Analysis of $L$-Stubborn Mining
We start by analysis of the revenue ratio of the adversary when it employs the $L$-stubborn mining strategy. Instead of building a complex Markov chain and going through the tedious analysis of steady-state distribution as it is done in [4], we can analyze the distinct possible paths an attack cycle can take.
@@
Proof:  The end of the unsuccessful attack cycle implies that the last block that arrives has to be honest and the adversary was never behind up until that the last block. As a result, the arrival sequence of honest and adversarial blocks creates a Dyck string with length $n<L$. There are $C_{n}$ number of such strings, where $C_{n}$ is the $n$-th Catalan number defined in (38) in Appendix A. Together with the last honest block each arrival sequence has probability $\alpha^{n}\beta^{n+1}$. Hence, an attack cycle ends unsuccessfully with the honest block at height $n+1$ with probability
@@
Proof:  First note that, in the limit, the longest chain is essentially the offset chain that grows whenever an adopt or override happens. As a result, the revenue ratio of the attacker is the number of the blocks created by the attacker in the offset chain divided by the total length of the offset chain. As the offset event is a stopping time, we can simply focus on a single attack cycle and calculate the expected number of blocks created by the adversary that become part of the offset chain denoted as $O_{A}$. Dividing this number to the expected growth of the offset chain in an attack cycle, denoted as $O_{T}$, gives us the revenue ratio. In other words, by strong law of large numbers, under the $L$-stubborn strategy, we have
$\rho_{L}=\lim_{t\to\infty}\frac{N_{A}^{t}}{N_{A}^{t}+N_{H}^{t}}=\frac{\mathbb{E}[O_{A}]}{\mathbb{E}[O_{T}]}.$ (16)
In Lemma 1 and Lemma 2, we provide the probability for each possible outcome of an attack cycle. If an attack cycle is successful, we have an $A$-chain at length $L$ and the maximum height of the honest block is $H=m<L$ at some time within the attack cycle. Clearly, if Lm>1L-m>1, the adversary will continue the cycle until the difference goes down to $1$. Thus, given the attack is successful, using Wald’s identity provided in Appendix A.3, the expected number of blocks that become part of the offset chain in this case is $L+\frac{(L-m-1)\alpha}{1-2\alpha}$ and all these blocks are adversarial. On the other hand, if the attack cycle is unsuccessful, $i$ adversarial blocks still make it into the offset chain w.p. $P_{u}(n,i)$ and the total number of blocks that make it into the offset chain is $n+1$. As a result, the expectation of the two cases are denoted as $L_{u,a}$ and $L_{u}$, respectively. Putting everything together, it is clear that $\mathbb{E}[O_{A}]=L_{s}+L_{u,a}$ and $\mathbb{E}[O_{T}]=L_{s}+L_{u}$, giving the desired result in (12). $\blacksquare$
@@ 4.1 Optimizing the Revenue Ratio
As we mentioned earlier, our $L$-stubborn mining strategy is a generalization and extension of the existing mining strategies of honest mining $(L=1)$, selfish mining [4] $(L=2)$, equal fork stubborn mining [6] $(L=\infty)$. As a result, for a given $(\alpha,\gamma)$ pair, a natural question is, what integer value of $L$ maximizes the revenue ratio $\rho_{L}$. To find the optimal $L$, one could evaluate $\rho_{L}$ in (12) and find the $L$ that provides the highest revenue ratio. As the calculations involve binomial coefficients and (12) involves fractions, it is not straightforward to find $L^{*}$. Instead, here, we provide a much simpler method to obtain the optimal $L$. To do so, we resort to the linearization method of the MDP problem [5] which is presented in Appendix B, and use it in a narrow-sense. Note that, every time a block is mined, the adversary faces a decision as to whether it should release $A$-chain and end the cycle, or continue. Clearly, if AH>1A-H>1, the decision should be to continue the attack cycle as the adversary can gain additional advantage at no-cost in these situations. On the other hand, if $A=H+1$, the risk is imminent as the adversary loses the advantage of overriding the $H$-chain if the next arrival is honest. As a result, we can narrow down the decision problem to a specific case where $A=H+1$ and analyze the situation therein via the linearization method.
@@
Next, using the definitions above and intermediary results in Appendix C, we provide the following quasiconcavity result related to the $L$-stubborn mining.
@@
Theorem 2 essentially says that $\rho_{L}$ is quasiconcave for integer $L$. We relegate the proof of Theorem 2 to the Appendix C as it requires some intermediary definitions and results. These results also shed light on a simple algorithm that can be used to find $L^{*}$.
Next, we provide Algorithm 1 that finds the optimal $L^{*}$ of $L$-stubborn mining. For that, let $\rho_{1}$, $\rho_{2}$ and $\rho_{\infty}$ be defined as in Definition 5, which are the revenue ratios of honest mining, selfish mining [4], and equal-fork stubborn mining [6, 37], respectively. Note that each of these values can be obtained by (12), where for $\rho_{\infty}$, $L_{s}$ can be set to zero as the $H$-chain will catch the $A$-chain eventually, i.e., no attack ends successfully, but the revenue comes from matching. In Appendix C, we use the intermediary results to show that Algorithm 1 outputs $L^{*}$, which essentially solves the decision problem mentioned earlier.
@@ 4.2 $(k+1)$-Stubborn Mining and Double-Spending
Note that, even when $L=k+1$, a successful attack cycle does not immediately imply a successful double-spending. It only implies that the $A$-chain will make it into the offset chain eventually. Similarly, an unsuccessful attack cycle does not immediately imply that $\mathsf{B}$ will be confirmed. Next, we provide the probability of each event regarding $\mathsf{B}$ and $\mathsf{B^{\prime}}$ for $(k+1)$-stubborn mining with $k$-confirmation rule.
@@
Proof:  For Double-Spending to happen, one possibility is, we will observe $A>k$ and $H\leq k$ at some point during the cycle. The first time this happens during the cycle is the first time $A=k+1$ and $H=m$ which happens w.p. $P_{s}(k+1,m)$. Until this point, the adversary matches every one of the $m$ honest blocks. From this point onwards, the adversary does not need to match the honest blocks as the strategy will eventually release the longer $A$-chain. Hence, $\mathsf{B}$ stays in the $H$-chain if none of the $m$ matching adversarial blocks make it into the prefix of the honest block at height $H$, which happens with probability $(1-\gamma)^{m-\mathbbm{1}_{m=k}}$. Note that, if $H=k$, then it does not matter whether prefix chain is after the $k$th block or not, hence $\mathbbm{1}_{m=k}$ at the exponent. Another possibility for Double-Spending is, if $H$-chain catches $A$-chain with both having length $k$ and the next block is honest which has probability $P_{u}(k)$. In this situation, if $H$-chain switches its prefix on the last honest block, which happens w.p. $(1-\gamma)^{k-1}\gamma$, then $\mathsf{B}$ gets discarded despite being confirmed earlier. For Service to happen, we will observe $L\geq H=A+1=m+1$, which ends the cycle with $H$-chain making it into the offset chain which happens w.p. $P_{u}(m)$. Further, $\mathsf{B}$ has to be in the prefix of the honest block at height $m+1$, which explains the term $(1-\gamma)^{m}$. $\blacksquare$
Here, we remind the reader that $\rho_{k+1}$ is the revenue ratio of the adversary when it employs the $(k+1)$-stubborn mining strategy and $\bar{L}=\sup\{L:\rho_{L}\geq\alpha\}$, i.e., the largest value of $L$ such that the revenue ratio of $L$-stubborn mining is higher than the honest mining strategy. Hence, due to the quasiconcavity proven in Theorem 2, under any $k$-confirmation rule with $k<\bar{L}$ is compromised by $(k+1)$-stubborn mining strategy since the adversary gains more revenue ratio by employing this strategy compared to the revenue ratio of the honest mining strategy, i.e., $\rho_{k+1}\geq\alpha$, and at the same time it can double-spend at no-cost. In other words, for a given $\alpha$ and $\gamma$, $k$-confirmation rule should be at least $k\geq\bar{L}$ to make sure that double-spending attack has loss of revenue. Note that, if $\rho_{\infty}\geq\alpha$, no $k$-confirmation is safe as double-spending has zero-cost. From this perspective, the safety of a blockchain system fails well before adversarial proportion reaches $\alpha$. We investigate this phenomenon further in the numerical results section by providing $L^{*}$ and $\bar{L}$ for each $\alpha$ and $\gamma$.
@@ 4.3 Combined Revenue Ratio
Proof:  Note that the first term in (18) provides the sum of successful cycles with double-spending events. Each term in that sum corresponds to the first time $A=k+1$ and $H=m$, and none of the matched adversarial blocks make it into the prefix of the honest block as explained in the proof of Theorem 3. Given Wald’s identity and the fact that a cycle will continue until the difference goes down to $1$, all but last $k$ adversarial blocks will result in double-spending events. Thus, total extra block rewards the adversary gets in a successful attack cycle in addition to the $(k+1)_{s}$ is provided as in (22), where the extra term is $\mathcal{R}$ times the expected number of honest blocks that are confirmed and later replaced in an attack cycle. An unsuccessful cycle can only result in double-spending if the event corresponding to the second term in (18) happens, where only one honest block is confirmed and replaced later, which explains the extra term in (23). $\blacksquare$
@@
Next, we provide another approach that incorporates the profit coming from double-spending when considering the $(k+1)$-stubborn mining attack provided here. As mentioned in [3], the value of service provided might be below the fee paid. In this case, we assume that each successful cycle results in only single double-spending event, i.e., only the first block $\mathsf{B}$ in the $H$-chain is attacked. We further assume that the value of service provided by the merchant equals $V$ and the adversary pays $F$ fee for the service with $F\geq V$, where both values are in block reward units. In this case, on average, the $(k+1)$-stubborn mining attack is more profitable than honest mining if